# On the spectral radius of the ratio of Girko matrices

### v1

@@ 2.3. Equilibrium measure
$\mathrm{d}X^{z}_{t}=-\frac{1}{2}X^{z}_{t}\mathrm{d}t+\sqrt{\frac{1+|z|^{2}}{n}}\mathrm{d}\mathcal{M}_{t},\qquad X^{z}_{0}=\frac{A-zB}{\sqrt{n}},$ (4.30)
where $\mathcal{M}_{t}$ is a standard complex Brownian motion matrix. For notational simplicity, we may often drop the dependence on $z$ or $t$ and write $X_{t}^{z}=(x_{ij})_{i,j\in[2n]}$. We use $X^{z}_{t}$ to construct the Hermitised matrix $H^{z}_{t}=(h_{ij})_{i,j\in[2n]}$ as in (4.7) and define the corresponding Green function $G^{z}_{t}(w)=(H^{z}_{t}-w)^{-1}$ as in (4.8). In particular,
$h_{aB}=\overline{h_{Ba}}=x_{a,B-n},\quad\qquad h_{Ba}=\overline{h_{aB}},\quad\qquad a\in[n],~B\in[n,2n],$ (4.31)
and all the other components of $H^{z}_{t}$ are equal to zero. Note that $(x_{ij})_{i,j\in[n]}$ are time-dependent random variables which also depend on $z$. Recall that the $(p,q)$-cumulant of a complex-valued random variable $\chi$ with finite moments are defined to be
@@
From (C2)-(C3), it is easy to check that second cumulants of $(x_{ij})_{i,j\in[n]}$ are invariant with time
$\kappa^{(1,1)}(x_{ij})=\frac{1+|z|^{2}}{n},\qquad\kappa^{(0,2)}(x_{ij})=\kappa^{(2,0)}(x_{ij})=0,$ (4.33)
@@
$|\kappa^{(p,q)}(x_{ij})|\lesssim n^{-2-c_{0}},\qquad p+q=3,4,\qquad|\kappa^{(p,q)}(x_{ij})|\lesssim n^{-\frac{p+q}{2}},\qquad p+q\geq 5,$ (4.34)
uniformly for any $t\geq 0$ and $|z|\leq C_{0}$.
Recall the definition of $I^{\geq\eta_{0}}$ in (4.13). Note that $\mathrm{Tr}G^{z}(\mathrm{i}\eta)$ is purely imaginary due to the block structure of $H^{z}$ in (4.7), so we have $\mathrm{Tr}G^{z}(\mathrm{i}\eta)=\mathrm{i}\operatorname{Im}\mathrm{Tr}G^{z}(\mathrm{i}\eta)$. Then we define the following short-hand
with $T=n^{100}$ and $\eta_{0}=n^{-1-\epsilon}$ for $0<\epsilon<c_{0}/100$.
We then use Itô’s formula to derive the dynamics of $F(I^{\geq\eta_{0}}_{t})$ as a function of $X^{z}_{t}=\{x_{ij}\}_{i,j\in[n]}$ given by (4.30). By direct computations and the relation (4.31), we obtain
$\frac{\mathrm{d}F(I^{\geq\eta_{0}}_{t})}{\mathrm{d}t}=\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\left(-\frac{1}{2}h_{aB}\frac{\partial F(I^{\geq\eta_{0}}_{t})}{\partial h_{aB}}-\frac{1}{2}\overline{h_{aB}}\frac{\partial F(I^{\geq\eta_{0}}_{t})}{\partial\overline{h_{aB}}}+\frac{1+|z|^{2}}{n}\frac{\partial^{2}F(I^{\geq\eta_{0}}_{t})}{\partial h_{aB}\partial\overline{h_{aB}}}\right)\mathrm{d}t+\mathrm{d}M,$ (4.36)
where $\mathrm{d}M$ denotes the corresponding diffusion term with vanishing expectation, due to the martingale property of Brownian motions. We next take the expectation of the remaining drift terms in (4.36). Note that the differentiation rules below are obtained from the definition of resolvent in (4.8):
$\frac{\partial(G^{z}_{t})_{xy}}{\partial h_{aB}}=-(G^{z}_{t})_{xa}(G^{z}_{t})_{By},\qquad\frac{\partial(G^{z}_{t})_{xy}}{\partial\overline{h_{aB}}}=-(G^{z}_{t})_{xB}(G^{z}_{t})_{ay},$ (4.37)
for any $x,y\in[2n]$, $a\in[n]$, and $B\in[n,2n]$. Recall that $\kappa^{(p+1,q)}(h_{aB})$ and $\kappa^{(p+1,q)}(\overline{h_{aB}})$ are the $(p+1,q)$-cumulants of the complex valued random variables from (4.31). Applying the cumulant expansion formula (for a formal statement, see (He and Knowles, 2017, Lemma 7.1)) to the expectation of the first two terms in (4.36), the resulting second-order cumulant terms from (4.33) will cancel precisely with the last drift term in (4.36). This key cancellation is a consequence of the fact that the second moments of the matrix entries of $X^{z}_{t}$ do not change with time. Hence, we start with the third-order terms ($p+q+1=3$) in the cumulant expansions, i.e.,
$\frac{\mathrm{d}\mathbb{E}\big[F(I^{\geq\eta_{0}}_{t})\big]}{\mathrm{d}t}=$ $-\frac{1}{2}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\left(\sum_{p+q+1=3}^{K_{0}}\frac{\kappa^{(p+1,q)}(h_{aB})}{p!q!}\mathbb{E}\left[\frac{\partial^{p+q+1}F(I^{\geq\eta_{0}}_{t})}{\partial h_{aB}^{p+1}\partial\overline{h_{aB}}^{q}}\right]\right)$
@@
$-\frac{1}{2}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\left(\sum_{p+q+1=3}^{K_{0}}\frac{\kappa^{(p+1,q)}(\overline{h_{aB}})}{p!q!}\mathbb{E}\left[\frac{\partial^{p+q+1}F(I^{\geq\eta_{0}}_{t})}{\partial\overline{h_{aB}}^{p+1}\partial h_{aB}^{q}}\right]\right)+O(n^{-10}),$ (4.38)
@@
where we stopped the expansion at a sufficiently high order with $p+q+1=K_{0}$ such that the truncation error is bounded by $O(n^{-10})$, using the cumulant bounds in (4.34), the differentiation rule in (4.37), the deterministic bound $\|G_{t}^{z}(\mathrm{i}\eta)\|\leq\eta^{-1}$, together with the local law estimates in (4.48) below. Such truncation argument is standard and frequently used in previous works e.g., He and Knowles (2017); Schnelli and Xu (2022); Cipolloni et al. (2024b), so we omit the details. We remark that the last error term holds uniformly for any $|z|\leq C_{0}$ and $0\leq t\leq 100\log n$.
@@
To estimate the third and higher order terms with $3\leq p+q+1\leq K_{0}$, we introduce the following lemma, whose proof is postponed till the end of this section.
@@
$\sup_{t\in[0,t_{0}]}\Big|\frac{\partial^{k}I^{\geq\eta_{0}}_{t}}{\partial(h_{aB})^{k_{1}}(\overline{h_{aB}})^{k_{2}}}\Big|\prec n^{2k\epsilon+\nu},\qquad t_{0}=100\log n.$ (4.39)
@@
Note that $H^{z}_{t}$ is defined as in (4.7) with the time dependent matrix $X^{z}_{t}$ from (4.30), which in distribution equals to
$X^{z}_{t}\stackrel{{\scriptstyle{\rm d}}}{{=}}\mathrm{e}^{-\frac{t}{2}}X^{z}+\sqrt{1-\mathrm{e}^{-t}}X^{\mathrm{Gin}},\qquad t\geq 0,$
with $X^{z}_{0}=X^{z}=(A-zB)/\sqrt{n}$ and $X^{z}_{\infty}=X^{\mathrm{Gin}}$ being the Gaussian counterpart which is independent of $X^{z}$. Using the resolvent identity, we obtain
@@
where we also used that $\|G(\mathrm{i}\eta)\|\leq\eta^{-1}$ and $|x_{ij}|\prec n^{-1/2}$ from condition (C3). Thus using that $F$ has uniformly bounded derivative and $\int|\Delta_{z}f(z)|\mathrm{d}^{2}z\leq n^{\nu}$ with $\nu<10^{-4}c_{0}$, we obtain from (4.35) that
@@
Recall $I^{\geq\eta_{0}}_{t}$ in (4.35). Using (4.37) and $G^{2}=\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}\eta}G(\mathrm{i}\eta)$, we further have
$\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h_{aB}}=$ $-\frac{\mathrm{i}}{4\pi}\int\Delta_{z}f_{n}(z)\int_{\eta_{0}}^{T}\big[\big(G_{t}^{z}(\mathrm{i}\eta)\big)^{2}\big]_{Ba}\mathrm{d}\eta\mathrm{d}^{2}z=\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{Ba}\mathrm{d}^{2}z+O_{\prec}(n^{-10}),$ (4.45)
@@
$\max_{x,y\in[2n]}\Big|\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{xy}\Big|\prec n^{2\epsilon},\qquad\mbox{~with~}\quad\eta_{0}=n^{-1-\epsilon}.$
Due to the Lipschitz continuity of $G^{z}_{t}$ in both $t$ and $z$, one can show that the above bound indeed holds true simultaneously for any $t\in[0,t_{0}]$ with $t_{0}=100\log n$ and $|z|\leq C_{0}$, i.e.,
$\sup_{t\in[0,t_{0}]}\sup_{|z|\leq C_{0}}\left\{\max_{x,y\in[2n]}\Big|\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{xy}\Big|\right\}\prec n^{2\epsilon},\qquad\mbox{~with~}\quad\eta_{0}=n^{-1-\epsilon},$ (4.48)
using a grid argument and union bound. Hence using that $\int|\Delta_{z}f_{n}(z)|\mathrm{d}^{2}z\leq n^{\nu}$, we conclude from (4.45) that
$\sup_{t\in[0,t_{0}]}\Big|\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h_{aB}}\Big|\prec n^{2\epsilon+\nu}.$ (4.49)
The same result also applies to $\partial/\partial\overline{h_{aB}}$. In general, for any $k\geq 1$, using the differentiation rules in (4.37) and the local law estimates in (4.48) repeatedly, we obtain the estimate for any $k$-th derivatives of $I^{\geq\eta_{0}}_{t}$ as stated in (4.39). This ends the proof of Lemma 4.4. ∎
@@ 5. Lemmas on weak convergence
We collect here classical results used in the proof of Theorem 1.2. The first one states that we can consider compactly supported smooth functions instead of continuous ones in the definition of weak convergence. This is a natural and familiar variant of the Portmanteau theorem that we were not able to locate in a textbook. The second one states that the convergence of point processes is characterized by compactly supported smooth test functions. The third and final one is about the delta method in the context of point process.
@@
Next, assuming that (5.1) is true for compactly supported smooth functions, we may show that $(X_{n})_{n\geq 1}$ is tight. For this, we may take a smooth function $g:\mathbb{R}\to[0,1]$ supported on [1,1][-1,1] such that $g|_{[-1/2,1/2]}\equiv 1$. Using that $\lim_{M\to\infty}g(x/M)=1$ and that $g$ is bounded by $1$, we obtain that $\lim_{M\to\infty}\mathbb{E}[g(X/M)]=1$. Then, for any $\varepsilon>0$, we may find an $M>0$ such that $\mathbb{E}[f(X/M)]>1-\varepsilon$. Finally, since $\lim_{n\to\infty}\mathbb{E}[g(X_{n}/M)]=\mathbb{E}[g(X/M)]$, we know that $\mathbb{E}[g(X_{n}/M)]>1-\varepsilon$ for $n$ large enough so that
@@
$\int_{\mathbb{C}}f\mathrm{d}\mu_{\mathcal{X}_{n}}\xrightarrow[n\to\infty]{\mathrm{law}}\int_{\mathbb{C}}f\mathrm{d}\mu_{\mathcal{X}}$
@@
$CR\varepsilon_{n}\sum_{y\in\mathcal{Y}_{n}}g(\sqrt{n}y)\xrightarrow[n\to\infty]{\mathrm{law}}0.$

### v2

@@ 2.3. Equilibrium measure
$\mathrm{d}M_{t}=$ $\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\sum_{l=1}^{2}\left(\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h^{(l)}_{aB}}\frac{\mathrm{d}\mathcal{M}^{(l)}_{aB}}{\sqrt{n}}+\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial\overline{h^{(l)}_{aB}}}\frac{\big(\mathrm{d}\mathcal{M}^{(l)}_{aB}\big)^{*}}{\sqrt{n}}\right).$ (4.35)
Note that the differentiation rules below are obtained from the definition of resolvent in (4.30) and (4.31):
$\frac{\partial(G^{z}_{t})_{ij}}{\partial h^{(1)}_{aB}}=-(G^{z}_{t})_{ia}(G^{z}_{t})_{Bj},\qquad\frac{\partial(G^{z}_{t})_{ij}}{\partial\overline{h^{(1)}_{aB}}}=-(G^{z}_{t})_{iB}(G^{z}_{t})_{aj},$ (4.36)
$\frac{\partial(G^{z}_{t})_{ij}}{\partial h^{(2)}_{aB}}=z(G^{z}_{t})_{ia}(G^{z}_{t})_{Bj},\qquad\frac{\partial(G^{z}_{t})_{ij}}{\partial\overline{h^{(2)}_{aB}}}=\overline{z}(G^{z}_{t})_{iB}(G^{z}_{t})_{aj},$ (4.37)
for any $i,j\in[[2n]]$, $a\in[[n]]$, and $B\in[[n+1,2n]]$.
Since $F$ is continuously differentiable up to fifth order, by Itô’s formula again and taking the expectation, we have
$\mathrm{d}\mathbb{E}\big[F(I^{\geq\eta_{0}}_{t})\big]=\mathbb{E}\Big[F^{\prime}(I^{\geq\eta_{0}}_{t})\big(\Theta_{1}+\Theta_{2}+\Theta_{3}\big)\Big]\mathrm{d}t+\frac{1}{2}\mathbb{E}\Big[F^{\prime\prime}(I^{\geq\eta_{0}}_{t})\big\langle\mathrm{d}M_{t},\mathrm{d}M_{t}\big\rangle\Big],$ (4.38)
using the martingale property of Brownian motions. By direct computations via (4.35), we have
$\frac{1}{2}\mathbb{E}\Big[F^{\prime\prime}(I^{\geq\eta_{0}}_{t})\big\langle\mathrm{d}M_{t},\mathrm{d}M_{t}\big\rangle\Big]=$ $\frac{1}{n}\sum_{l=1}^{2}\mathbb{E}\left[F^{\prime\prime}(I^{\geq\eta_{0}}_{t})\left(\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h^{(l)}_{aB}}\right)\left(\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial\overline{h^{(l)}_{aB}}}\right)\right]\times\mathrm{d}t,$ (4.39)
where we also used the independence between $\mathcal{M}^{(1)}_{t}$ and $\mathcal{M}^{(2)}_{t}$. To compute the first drift term in (4.38), we apply the cumulant expansion formula (for a formal statement, see (He and Knowles, 2017, Lemma 7.1)) with respect to $\big(h^{(l)}_{aB}\big)$. Recall that the $(p,q)$-cumulant of a complex-valued random variable $\chi$ with finite moments are defined to be
@@
From (C2)-(C3), it is easy to check that second cumulants of $\big(h^{(l)}_{aB}\big)$ are invariant with time, i.e.,
$\kappa^{(1,1)}\big(h^{(l)}_{aB}\big)=\frac{1}{n},\qquad\kappa^{(0,2)}\big(h^{(l)}_{aB}\big)=\kappa^{(2,0)}\big(h^{(l)}_{aB}\big)=0,$ (4.41)
@@
$|\kappa^{(p,q)}\big(h^{(l)}_{aB}\big)|\lesssim n^{-2-c_{0}},\qquad p+q=3,4,\qquad|\kappa^{(p,q)}\big(h^{(l)}_{aB}\big)|\lesssim n^{-\frac{p+q}{2}},\qquad p+q\geq 5,$ (4.42)
uniformly for any $t\geq 0$. We then write out the second order terms in the expansions, i.e.,
$\mathbb{E}\big[F^{\prime}(I^{\geq\eta_{0}}_{t})\Theta_{1}\big]=$ $-\frac{1}{2}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\sum_{l=1}^{2}\mathbb{E}\left[h^{(l)}_{aB}F^{\prime}(I^{\geq\eta_{0}}_{t})\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h^{(l)}_{aB}}\right$]
@@
$=$ $-\frac{1}{2n}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\sum_{l=1}^{2}\mathbb{E}\left[\frac{\partial}{\partial\overline{h^{(l)}_{aB}}}\left\{F^{\prime}(I^{\geq\eta_{0}}_{t})\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h^{(l)}_{aB}}\right\}\right]+\mbox{higher order terms},$ (4.43)
using (4.41). The second drift term in (4.38) with $\Theta_{2}$ can be handled very similarly with $h^{(l)}_{aB}$ replaced by $\overline{h^{(l)}_{aB}}$. Then we observe that the above second order cumulant expansion terms will cancel precisely with the third drift term in (4.34) and the quadratic variation term in (4.39). This key cancellation is a consequence of the fact that the second moments of the matrix entries in (4.29) do not change with time. Therefore, the remaining terms in (4.38) are the higher order cumulant expansion terms in (4.43). More precisely, the cumulant expansions starting with the third-order terms ($p+q+1=3$) are given by
$\frac{\mathrm{d}\mathbb{E}\big[F(I^{\geq\eta_{0}}_{t})\big]}{\mathrm{d}t}=$ $-\frac{1}{2}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\sum_{l=1}^{2}\left(\sum_{p+q+1=3}^{5}\frac{\kappa^{(p+1,q)}(h^{(l)}_{aB})}{p!q!}\mathbb{E}\left[\frac{\partial^{p+q+1}F(I^{\geq\eta_{0}}_{t})}{\partial(h^{(l)}_{aB})^{p+1}\partial(\overline{h^{(l)}_{aB}})^{q}}\right]\right)$
$-\frac{1}{2}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\sum_{l=1}^{2}\left(\sum_{p+q+1=3}^{4}\frac{\kappa^{(p+1,q)}(\overline{h^{(l)}_{aB}})}{p!q!}\mathbb{E}\left[\frac{\partial^{p+q+1}F(I^{\geq\eta_{0}}_{t})}{\partial(\overline{h^{(l)}_{aB}})^{p+1}\partial(h^{(l)}_{aB})^{q}}\right]\right)+O(n^{-1/2+c(\epsilon+\nu)}),$ (4.44)
@@
where $\kappa^{(p+1,q)}(\cdot)$ are the cumulants as bounded in (4.42) and we stopped the expansions at the fourth order with $p+q+1=4$ such that the truncation error is bounded by $O(n^{-1/2+c(\epsilon+\nu)})$ for some constant $c>0$. To estimate the truncation error, we used the cumulant bounds in (4.42), the differentiation rules in (4.36)-(4.37), the deterministic bound $\|G_{t}^{z}(\mathrm{i}\eta)\|\leq\eta^{-1}$, together with the estimates in Lemma 4.4 below. Such truncation argument is standard and frequently used in previous works e.g., He and Knowles (2017); Schnelli and Xu (2022); Cipolloni et al. (2024b), so we omit the details. We remark that the last error term holds uniformly for any $0\leq t\leq 100\log n$.
@@
To estimate the third and higher order terms in (4.1), we introduce the following lemma, whose proof is postponed till the end of this section.
@@
$\sup_{t\in[0,t_{0}]}\left|\frac{\partial^{k}I^{\geq\eta_{0}}_{t}}{\partial(h^{(l)}_{aB})^{k_{1}}(\overline{h^{(l)}_{aB}})^{k_{2}}}\right|\prec n^{2k\epsilon+\nu},\qquad l=1,2,\qquad t_{0}=100\log n.$ (4.45)
@@
Note that $H^{z}_{t}$ is defined as in (4.30) with the time dependent matrix $X^{z}_{t}={\tilde{A}}_{t}-z{\tilde{B}}_{t}$ from (4.29). For any fixed $t\geq 0$, $X^{z}_{t}$ in distribution equals to
$X^{z}_{t}\stackrel{{\scriptstyle{\rm d}}}{{=}}\mathrm{e}^{-\frac{t}{2}}X^{z}+\sqrt{1-\mathrm{e}^{-t}}X^{\mathrm{Gin}},$
with $X^{z}_{0}=X^{z}=(A-zB)/\sqrt{n}$ and $X^{z}_{\infty}=(A^{\mathrm{Gin}}-zB^{\mathrm{Gin}})/\sqrt{n}$ being the Ginibre counterpart which is independent of $X^{z}$. Using the resolvent identity, we obtain
@@
where we also used that $\|G(\mathrm{i}\eta)\|\leq\eta^{-1}$ and $|x_{ij}|\prec n^{-1/2}$ from condition (C3). Thus using that $F$ has uniformly bounded derivatives up to fifth order and $\int|\Delta_{z}f(z)|\mathrm{d}^{2}z\leq n^{\nu}$ with $\nu<10^{-4}c_{0}$, we obtain from (4.32) that
@@
Recall $I^{\geq\eta_{0}}_{t}$ in (4.32) and the differentiation rules in (4.36)-(4.37). Below we will present the proof for $l=1$ using (4.36), and the proof for $l=2$ is very similar using (4.37) and that $f_{n}(z)=0$ for $|z|>C_{0}$. Using (4.36) and $G^{2}=\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}\eta}G(\mathrm{i}\eta)$, we have
$\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h^{(l)}_{aB}}$= $-\frac{\mathrm{i}}{4\pi}\int\Delta_{z}f_{n}(z)\int_{\eta_{0}}^{T}\big[\big(G_{t}^{z}(\mathrm{i}\eta)\big)^{2}\big]_{Ba}\mathrm{d}\eta\mathrm{d}^{2}z=\frac{1}{4\pi}\int\Delta_{z}f_{n}(z)\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{Ba}\mathrm{d}^{2}z+O_{\prec}(n^{-10}),$ (4.51)
@@
$\max_{x,y\in[[2n]]}\Big|\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{xy}\Big|\prec n^{2\epsilon},\qquad\mbox{~with~}\quad\eta_{0}=n^{-1-\epsilon}.$
Due to the Hölder continuity of $G^{z}_{t}$ in both $t$ and $z$, one can show that the above bound indeed holds true simultaneously for any $t\in[0,t_{0}]$ with $t_{0}=100\log n$ and $|z|\leq C_{0}$, i.e.,
$\sup_{t\in[0,t_{0}]}\sup_{|z|\leq C_{0}}\left\{\max_{x,y\in[[2n]]}\Big|\big(G^{z}_{t}(\mathrm{i}\eta_{0})\big)_{xy}\Big|\right\}\prec n^{2\epsilon},\qquad\mbox{~with~}\quad\eta_{0}=n^{-1-\epsilon},$ (4.54)
using a grid argument and union bound (see similar explanations below (4.17)). Hence using that $\int|\Delta_{z}f_{n}(z)|\mathrm{d}^{2}z\leq n^{\nu}$, we conclude from (4.51) that
$\sup_{t\in[0,t_{0}]}\Big|\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h^{(l)}_{aB}}\Big|\prec n^{2\epsilon+\nu}.$ (4.55)
The same result also applies to $\partial/\partial\overline{h^{(l)}_{aB}}$. In general, for any $k\geq 1$, using the differentiation rules in (4.36)-(4.37) and the local law estimates in (4.54) repeatedly, we obtain the estimate for any $k$-th derivatives of $I^{\geq\eta_{0}}_{t}$ as stated in (4.45). This ends the proof of Lemma 4.4. ∎
@@ 5. Lemmas on weak convergence
We collect here classical results used in the proof of Theorem 1.2 that we were not able to locate in a textbook, and we provide abridged proofs for the convenience of readers.
The first one states that we can consider compactly supported smooth functions instead of continuous ones in the definition of weak convergence. This is a natural and familiar variant of the Portmanteau theorem. The second one states that the convergence of point processes is characterized by compactly supported smooth test functions. The third and final one is about the delta method in the context of point process.
@@
Next, assuming that (5.1) is true for compactly supported smooth functions, we may show that $(X_{n})_{n\geq 1}$ is tight. For this, we may take a smooth function $g:\mathbb{R}\to[0,1]$ supported on [1,1][-1,1] such that $g|_{[-1/2,1/2]}\equiv 1$. Using that $\lim_{M\to\infty}g(x/M)=1$ and that $g$ is bounded by $1$, we obtain that $\lim_{M\to\infty}\mathbb{E}[g(X/M)]=1$. Then, for any $\varepsilon>0$, we may find an $M>0$ such that $\mathbb{E}[g(X/M)]>1-\varepsilon$. Finally, since $\lim_{n\to\infty}\mathbb{E}[g(X_{n}/M)]=\mathbb{E}[g(X/M)]$, we know that $\mathbb{E}[g(X_{n}/M)]>1-\varepsilon$ for $n$ large enough so that
@@
$\int_{\mathbb{C}}f\mathrm{d}\mu_{\mathcal{X}_{n}}\xrightarrow[n\to\infty]{\mathrm{d}}\int_{\mathbb{C}}f\mathrm{d}\mu_{\mathcal{X}}$
@@
$CR\varepsilon_{n}\sum_{y\in\mathcal{Y}_{n}}g(\sqrt{n}y)\xrightarrow[n\to\infty]{\mathrm{d}}0.$